please help me solve 2x/7 - 3x/5 = 6/35
please explain how you get
answer

Found 2 solutions by checkley75, rmromero:
Answer by checkley75(3666)

2X/7-3X/5=6/35
(5*2X-7*3X)/7*5=6/35
(10X-21X)/35=6/35 BECAUSE THERE ARE COMMON DENOMINATORS THEY CAN BE CANCELLED OUT THUS WE HAVE:
-11X=6
X=6/-11
X=-6/11 ANSWER.
2(-6/11)/7-3(-6/11)/5=6/35
(-12/11)/7+(18/11)/5=6/35
(5*-12/11+7*18/11)/35=6/35
(-60/11+126/11)=6
66/11=6
6=6

Answer by rmromero(383)


please help me solve 2x/7 - 3x/5 = 6/35
please explain how you get
answer
2x 3x 6
____ - _____ = _____
7 5 35
We must get rid of the denominators so that we can solve for x.
We will do that by multiplying LCD each term.
LCD = 35
2x 3x 6
(35) ____ - _____ (35 )= _____ (35)
7 5 35
first term: 35/7 = 5. Multiply it to 2x
Second term: 35/5= 7. Multiply it to 3x. Copy the sign
Third Term: 35/35 = 1. Multiply it to 6
10x - 21x = 6
- 11x = 6 divide both sides by -11
-6
x = ____
11
Checking:
2x 3x 6
____ - _____ = _____
7 5 35


2(-6/11) 3(-6/11) 6
_________ - _________ = _____
7 5 35


-12/11 -18/11 6
_________ - _________ = _____
7 5 35

-12/11 -18/11 6
(35)_________ - _________ (35) = _____ (35)
7 5 35

-12 -18
______ * 5 - ______ * 7 = 6
11 11

-60 -126
______ - ______ = 6
11 11
-60 126
______ + ______ = 6
11 11
66
____ = 6
11
6 = 6 --------> True